OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

hartnn (hartnn):

use this slope point formula y-y1 = m (x-x1) where m =- 2 x1 = 5 y1 = -1 are given

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is is B?

OpenStudy (anonymous):

is it i mean

hartnn (hartnn):

no.... y- y1 -----> y - (-1) = y+1 on the left side

OpenStudy (anonymous):

lol i suck at moth sorry

OpenStudy (anonymous):

math

hartnn (hartnn):

practice will make you good in math :) keep practicing!

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^